# ViewMask-1-to-3: Multi-View Consistent Image Generation via Multimodal Diffusion Models

### v1

@@ Abstract
Multi-view image generation from a single image and text description remains challenging due to the difficulty of maintaining geometric consistency across different viewpoints. Existing approaches typically rely on 3D-aware architectures or specialized diffusion models that require extensive multi-view training data and complex geometric priors. In this work, we introduce ViewMask-1-to-3, a pioneering approach to apply discrete diffusion models to multi-view image generation. Unlike continuous diffusion methods that operate in latent spaces, ViewMask-1-to-3 formulates multi-view synthesis as a discrete sequence modeling problem, where each viewpoint is represented as visual tokens obtained through MAGVIT-v2 tokenization. By unifying language and vision through masked token prediction, our approach enables progressive generation of multiple viewpoints through iterative token unmasking with text input. ViewMask-1-to-3 achieves cross-view consistency through simple random masking combined with self-attention, eliminating the requirement for complex 3D geometric constraints or specialized attention architectures. Our approach demonstrates that discrete diffusion provides a viable and simple alternative to existing multi-view generation methods, ranking first on average across GSO and 3D-FUTURE datasets in terms of PSNR, SSIM, and LPIPS, while maintaining architectural simplicity.
Figure 1: Comparison of multi-view image generation approaches. (Left-Top) 3D representation methods require expensive 3D reconstruction and rendering. (Left-Bottom) Camera-conditioned diffusion methods need precise camera parameters. (Left-middle) Image editing models generate views sequentially without consistency guarantees. (Right) Our discrete diffusion approach enables parallel multi-view generation through unified sequence modeling with flexible text-based control.
@@ 1 Introduction
Multi-view image generation has emerged as a fundamental challenge [he2024head360, liu2024mvsgaussian, he2024emotalk3d], with applications spanning virtual reality [zhou2025cross] and 3D reconstruction [alzayer2024seeing]. The task aims to generate multiple consistent viewpoints of an object or scene from limited input—typically a single image and accompanying text description. Computational approaches face significant challenges in maintaining geometric consistency, preserving fine-grained details, and ensuring semantic coherence across viewpoints.
Traditional multi-view synthesis approaches primarily rely on explicit 3D representations or geometry-aware neural architectures. 3D-aware GANs [chan2021pi, schwarz2020graf] leverage neural radiance fields or voxel grids to enforce geometric constraints, while recent diffusion-based methods [shi2024mvdream, liu2024syncdreamer] employ specialized architectures with view-dependent conditioning. However, these approaches often require extensive multi-view training datasets, complex geometric priors, or computationally expensive 3D intermediate representations. Moreover, they typically operate in continuous latent spaces, limiting precise controllability over the generation process.
Recent advances in discrete diffusion models [nie2025large, ye2025dream, sahoo2024simple, you2025lladav] have demonstrated remarkable success in language modeling and multimodal understanding by treating generation as progressive token unmasking, models like LLaDA [nie2025large] for language generation, LLaDA-V [you2025lladav] for vision-language understanding, MMaDA [yang2025mmada] and Lumina-DiMOO [xin2025luminadimooomnidiffusionlarge] for multimodal generation, enable parallel decoding and bidirectional context modeling. These models operate by progressively revealing masked tokens in a learned sequence, enabling iterative refinement and controllable generation.
Discrete diffusion models offer two significant benefits: (1) Unified architecture across modalities, aligning with the industry shift toward integrated frameworks and enabling seamless cross-modal integration [huunified2023udd, yang2025mmada]. (2) Efficient parallel processing that surpasses sequential autoregressive methods in both speed and flexibility through simultaneous token prediction. The inherent bidirectional attention naturally captures contextual relationships, making discrete diffusion uniquely suited for tasks requiring consistency across complex dependencies [chang2022maskgit, xin2025luminadimooomnidiffusionlarge].
@@
In this work, we introduce ViewMask-1-to-3, a novel approach that adapts discrete diffusion models to multi-view image generation. Our key insight is that multi-view generation can be naturally formulated as a sequence modeling problem, where each viewpoint is represented as a sequence of discrete visual tokens obtained through MAGVIT-v2 [yu2024language] tokenization. By extending masked token prediction from language to the vision, ViewMask-1-to-3 enables progressive generation of multiple viewpoints through iterative token unmasking.
Unlike existing approaches [liu2023zero, kwak2023vivid, shi2023zero123++, zhang2025ar] that rely on precise camera parameters [liu2023zero, zhang2025ar], 3D reconstruction rendering [tangdreamgaussian, liu2024syncdreamer], or large image editing models that can only generate a single view at once with coarse view control, ViewMask-1-to-3 achieves cross-view consistency through simple random masking combined with bidirectional attention, as illustrated in Figure 1. This approach allows for parallel refinement of visual tokens across all viewpoints, while the iterative nature of masked diffusion provides opportunities for progressive quality improvement during the generation process. Our contributions are as follows:
  • We present a novel discrete diffusion method for multi-view image generation task, demonstrating how masked token prediction can effectively address visual synthesis across multiple viewpoints.
@@
  • The proposed simple masking strategy, when combined with bidirectional attention mechanisms, naturally achieves cross-view consistency without requiring complex 3D geometric priors or specialized architectures.
  • We evaluate ViewMask-1-to-3 on standard multi-view generation benchmarks, achieving the best average performance on both GSO and 3D-FUTURE datasets, establishing discrete diffusion as a viable alternative to existing continuous diffusion and 3D-aware methods.
@@ 2.1 Multi-View Image Generation
Early approaches ensure geometric consistency through explicit 3D modeling. 3D-aware GANs such as $\pi$-GAN [chan2021pi] and EG3D [chan2022efficient] leverage neural radiance fields or tri-plane representations for view-consistent synthesis. More recently, DreamGaussian [tangdreamgaussian] employs 3D Gaussian Splatting for efficient content creation; TRELLIS [xiang2024structured] adopts flow matching to obtain structured 3D latents and then extracts multi-view images; MVDream [shi2024mvdream] and SyncDreamer [liu2024syncdreamer] integrate NeRF [mildenhall2021nerf] into diffusion frameworks with synchronized denoising across viewpoints. While achieving strong geometric coherence, these methods require computationally expensive 3D representations and specialized rendering pipelines that increase system complexity.
@@
An alternative approach generates multi-view images directly in 2D space through camera-conditioned diffusion. Zero-1-to-3 family [liu2023zero, deitke2023objaverse, shi2023zero123++] pioneered fine-tuning pretrained diffusion models with camera pose conditioning for zero-shot novel view synthesis. Recent extensions include ViVid-1-to-3 [kwak2023vivid], which leverages video diffusion for temporal coherence; AR-1-to-3 [zhang2025ar], which introduces autoregressive viewpoint generation; EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], which enforces epipolar constraints through specialized attention mechanisms. However, these methods rely on explicit camera parameterization and specialized pose conditioning architectures, requiring precise viewpoint specifications and limiting flexibility in view specification through natural language.
@@
Recent methods attempt to generate multi-view images via image editing strategies. These approaches leverage various techniques: MV-Adapter [huang2024mvadapter] transforms text-to-image models into multi-view generators without altering their structure; ImageBrush [yang2023imagebrush] employs visual instructions rather than text for precise manipulation; SceneScape [fridman2023scenescape] ensures geometric consistency through depth-aware generation; and Lumina-DiMOO [xin2025luminadimooomnidiffusionlarge] introduces discrete diffusion in image editing with impressive results. While these methods rely on specialized editing strategies, ViewMask-1-to-3 naturally achieves cross-view consistency through token-level modeling using discrete diffusion, offering a more unified and elegant solution to multi-view synthesis.
@@ 2.2 Discrete Diffusion Models
D3PM [austin2021structured] established the foundational framework for discrete diffusion, extending traditional diffusion to discrete state spaces through structured denoising processes. DiffusionBERT [he2023diffusionbert] demonstrated that pre-trained BERT models could be adapted for text generation through discrete diffusion training. A significant breakthrough came with LLaDA [nie2025large], which showed that masked diffusion models can match autoregressive models like LLaMA3-8B in language generation while offering parallel decoding and bidirectional context modeling. LLaDA-V [you2025lladav] extended this framework to multimodal understanding, demonstrating the potential of discrete diffusion in vision-language tasks. These advances have established discrete diffusion as a viable alternative to autoregressive approaches, for tasks requiring controllable and parallel generation.
@@ 2.3 Visual Tokenization
The effectiveness of discrete diffusion for visual tasks critically depends on visual tokenization quality. VQVAE [oord2017neural] introduced discrete visual representations through vector quantization, laying the groundwork for discrete visual modeling. MaskGIT [chang2022maskgit] applied masked token prediction to image generation using vector-quantized representations. MAGVIT-v2 [yu2024language] enabling high-quality discrete tokenization for both images and videos through lookup-free quantization and architectural improvements. The success of these tokenizers has made discrete approaches increasingly competitive with continuous methods in visual generation tasks.
@@ 3.2 Model Design
We adopt MAGVIT-v2 [yu2024language] as our visual tokenizer due to its superior performance in discrete visual representation. For an input image $I_{i}$, the tokenizer produces a sequence of discrete visual tokens,
@@ 3.3 Multi-Stage Training Strategy
Stage1: Pretraing for Multimodal Alignment
@@
We construct a multimodal image–caption dataset by combining ImageNet [deng2009imagenet] images paired with captions from ImageNet-1K-VL-Enriched [VisualLayer2024ImageNetEnriched] and the LLaVA Visual Instruct CC3M Pretrain dataset [liu2023llava], which is a subset released by LLaVA from the CC-3M corpus, to facilitate cross-modal alignment.
@@
At this stage, the model is trained on a curated subset of Objaverse [deitke2023objaverse] and Habitat Synthetic Scene Dataset (HSSD) [khanna2023hssd]. Following the data split provided by Objaverse++ [lin2025objaversecurated3dobject], we removed samples labeled as Low Quality (No semantic meaning. Objects that annotators cannot identify or are corrupted), to ensure more stable training.
@@
We enhance Objaverse with Cap3D [luo2023scalable, luo2024view], which offers detailed textual descriptions for 3D assets.
@@ 4 Experiments
Our model was trained on a subset of Objaverse [deitke2023objaverse] and HSSD [khanna2023hssd]. Following the data split provided by Objaverse++ [lin2025objaversecurated3dobject], we removed samples labeled as Low Quality, to ensure more stable training. To enrich the textual supervision, we adopted the Cap3D [luo2023scalable, luo2024view] annotations, which provide descriptive captions for 3D objects. In total, we used 180K 3D objects, each rendered as an 8-frame RGBA orbit sequence at a resolution of 256×256, with the elevation fixed at 30°. During training, one frame was randomly selected as the reference image for each object, resulting in a multi-view generation dataset.
@@
We evaluate the multi-view generation results from two perspectives. First, for the directly generated novel views, we adopt common image-level metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [1284395], and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], to assess visual fidelity and perceptual quality. Then, we employ InstantMesh [xu2024instantmesh] to reconstruct 3D geometry from the generated views, and evaluate the resulting point clouds using Chamfer Distance (CD) and Intersection over Union (IoU) to quantify geometric consistency and accuracy.
Experimental Settings
[…]
@@ 4.1 Image to Multi-view Images
Method Architecture Control GSO 3D-FUTURE Rank
@@
Zero-1-to-3 [liu2023zero] 2D Cont. Diff. Camera Poses 18.8219 0.8294 0.1659 17.0526 0.8163 0.1760 5.3
Zero-1-to-3 XL [liu2023zero, objaverseXL] 2D Cont. Diff. Camera Poses 19.6839 0.8381 0.1518 18.4702 0.8337 0.1485 3.2
ViVid-1-to-3 [kwak2023vivid] 2D Cont. Diff. Camera Poses 19.7978 0.8566 0.1764 18.3241 0.8437 0.1682 3.5
Zero123++ [shi2023zero123++] 2D Cont. Diff. Fixed Views 19.6373 0.8045 0.3550 23.5001 0.8527 0.1529 4.3
Epidiff [huang2024epidiffenhancingmultiviewsynthesis] 2D Cont. Diff. Fixed Views 18.9917 0.8244 0.1667 17.4592 0.8108 0.1785 5.7
MV-Adapter [huang2024mvadapter] 2D Cont. Diff. Fixed Views 19.4673 0.7518 0.1503 19.3375 0.7217 0.4036 6.0
AR-1-to-3 [zhang2025ar] Video Cont. Diff. Fixed Views 13.5084 0.7376 0.3514 13.7724 0.7479 0.3034 8.5
TRELLIS [xiang2024structured] 3D Rectified Flow 3D Asset Render 14.4493 0.8058 0.3223 14.4604 0.7896 0.2990 7.2
@@
We compare our method against a comprehensive set of state-of-the-art diffusion-based novel view generation approaches, including Zero-1-to-3 [liu2023zero], Zero-1-to-3 XL [liu2023zero, objaverseXL], ViVid-1-to-3 [kwak2023vivid], Zero123++ [shi2023zero123++], MV-Adapter [huang2024mvadapter], EpiDiff [huang2024epidiffenhancingmultiviewsynthesis], AR-1-to-3 [zhang2025ar], and TRELLIS [xiang2024structured].
@@
To quantitatively evaluate the zero-shot generalization capability of our model, we conduct experiments on two aspects: (1) image-level quality, measured on novel views generated from a single input image, and (2) 3D-level geometric accuracy, evaluated on 3D reconstructions. These experiments are evaluated on the Google Scanned Objects (GSO) dataset [xin2025luminadimooomnidiffusionlarge] and the 3D-FUTURE dataset [fu20203dfuture3dfurnitureshape].
The input views are selected with controlled offsets in azimuth and elevation relative to the canonical frontal view of each 3D object. To account for inconsistent pose assumptions across baselines, we render ground-truth multi-view images using each model’s default viewing configuration. For models that support arbitrary-angle rendering, such as Zero-1-to-3 and Zero-1-to-3 XL, the views are rendered at the same angles as those used by ViewMask-1-to-3 to ensure a fair comparison. We then evaluate PSNR, SSIM, and LPIPS following each model’s respective settings, and report the results in Tab. 1. Overall, our method, ViewMask-1-to-3, achieves the best combined performance across both datasets and attains the highest rank (1.7), demonstrating its superiority in multi-view generation.
To evaluate geometric consistency, we reconstruct 3D shapes from the generated multi-view images for each model. We then compute CD and IoU between the reconstructed shapes and the ground-truth meshes. The quantitative results are reported in Tab. 2, demonstrating that our method achieves lower CD and higher IoU, indicating more accurate 3D reconstruction and better cross-view consistency.
The performance of Zero123++ on the 3D-FUTURE dataset is lower than expected mainly due to its default generation of opaque images with a gray background (corresponding to the zero value in the Stable Diffusion VAE). Removing this background using rembg may inadvertently eliminate fine-grained structures in 3D-FUTURE objects, thereby degrading reconstruction quality. For TRELLIS, some generated 3D assets include an additional bottom platform, which reduces the effective object scale. The necessary post-processing for platform removal and alignment with the ground-truth (GT) geometry introduces additional geometric errors, leading to inferior quantitative results.

### v2

@@ Abstract
Motivated by discrete diffusion’s success in language-vision modeling, we explore its potential for multi-view generation, a task dominated by continuous approaches. We introduce ViewMask-1-to-3, formulating multi-view synthesis as a discrete sequence modeling problem where each viewpoint is represented as visual tokens from MAGVIT-v2. Through masked token prediction, our approach enables progressive multi-view generation via iterative token unmasking, unifying language and vision in a shared token space. Importantly, simple random masking combined with self-attention naturally encourages cross-view consistency without specialized architectures or 3D geometric priors. Our method outperforms the baseline on the GSO and 3D-FUTURE benchmarks, ranking first on average across standard image metrics and improving IoU by 10.6% on 3D-FUTURE. This validates discrete diffusion as a promising candidate for multi-view generation.
Figure 1: Comparison of multi-view image generation approaches. (Left-Top) 3D representation methods with 3D reconstruction and rendering. (Left-Bottom) Camera-conditioned diffusion methods with precise camera parameters. (Left-middle) Image editing models generate views with sequential manner. (Right) Our discrete diffusion approach enables parallel multi-view generation through unified sequence modeling with flexible text-based control.
@@ 1 Introduction
Multi-view image generation has emerged as a fundamental challenge [15, 29, 14], with applications spanning virtual reality [58] and 3D reconstruction [1]. The task aims to generate multiple consistent viewpoints of an object or scene from limited input—typically a single image and accompanying text description. Computational approaches face significant challenges in maintaining geometric consistency, preserving fine-grained details, and ensuring semantic coherence across viewpoints.
Recent advances in this domain have achieved impressive results through diverse technical approaches. 3D representation methods [46, 42, 30] leverage explicit 3D representations, such as neural radiance fields, 3D Gaussian primitives, or structured latent spaces, to ensure geometric consistency through native 3D modeling and rendering. Camera-conditioned diffusion models [28, 39, 22, 19] have demonstrated remarkable flexibility by directly generating multi-view images in 2D space, either controlling the views with precise camera parameters or directly producing a fixed views. Image Editing model [18, 31, 27] explore integrating multi-view generation with image editing manner, enabling seamless transitions between different vision tasks. These methods, predominantly built on continuous diffusion in latent space, have collectively established strong baselines.
Meanwhile, a distinct paradigm has shown remarkable success in unified language-vision modeling: discrete diffusion models based on masked token prediction [35, 54, 52]. By operating in a unified vocabulary space, they offer intriguing properties: (1) native multi-modal fusion of text and visual tokens, (2) bidirectional, parallel context modeling during generation, and (3) architectural alignment with large language models. This raises a compelling yet unexplored question: Can this discrete paradigm serve as a viable alternative for the geometrically demanding task of multi-view generation?
@@
In this work, we present ViewMask-1-to-3, an exploration of discrete diffusion models for multi-view consistent image generation. Our key insight is to formulate multi-view synthesis as a discrete sequence modeling problem, where each viewpoint is represented as a sequence of visual tokens obtained through learned tokenization [55]. By extending masked token prediction, a technique proven effective in language modeling [35], to the visual domain with multiple viewpoints, we enable progressive generation through iterative token unmasking. Importantly, we find that a simple random masking strategy, when combined with bidirectional self-attention, naturally encourages cross-view consistency without requiring specialized architectural components or explicit 3D geometric priors. Our contributions are as follows:
  • We present a systematic exploration of discrete diffusion models for multi-view image generation, demonstrating that masked token prediction can achieve competitive performance in this domain.
@@
  • The proposed simple masking strategy, when combined with bidirectional attention mechanisms, naturally encourages cross-view consistency without requiring 3D geometric priors or specialized architectures.
  • We evaluate ViewMask-1-to-3 on standard multi-view generation benchmarks, where it tops the average rankings on image-level metrics (PSNR, SSIM, LPIPS, CD, IOU) for GSO and 3D-FUTURE, particularly achieving a 10.6% higher IoU on 3D-FUTURE.
@@ 2.1 Multi-View Image Generation
Early approaches ensure geometric consistency through explicit 3D modeling. 3D-aware GANs such as $\pi$-GAN [5] and EG3D [4] leverage neural radiance fields or tri-plane representations for view-consistent synthesis. More recently, DreamGaussian [42] employs 3D Gaussian Splatting for efficient content creation; TRELLIS [46] adopts flow matching to obtain structured 3D latents and then extracts multi-view images; MVDream [40] and SyncDreamer [30] integrate NeRF [34] into diffusion frameworks with denoising across viewpoints.
@@
An alternative approach generates multi-view images directly in 2D space through camera-conditioned diffusion. Zero-1-to-3 [28] pioneered fine-tuning pretrained diffusion models with camera pose conditioning for zero-shot novel view synthesis, generating each view independently. Subsequent works address cross-view consistency through various strategies: Zero123++ [39] models the joint distribution of multiple views by concatenating them into a unified representation; SyncDreamer [30] and One-2-3-45 [27] incorporate synchronized denoising and post-processing refinement. Recent extensions include ViVid-1-to-3 [22], which leverages video diffusion for temporal coherence; MV-AR [17], which introduces autoregressive viewpoint generation; and EpiDiff [19], which enforces epipolar constraints through specialized attention mechanisms. However, these methods typically focus on image-to-multi-view synthesis, with text-to-multi-view requiring additional text-to-image models to generate reference images first.
@@
Recent methods attempt to generate multi-view images via image editing strategies. These approaches leverage various techniques: MV-Adapter [18] transforms text-to-image models into multi-view generators without altering their structure; ImageBrush [53] employs visual instructions rather than text for precise manipulation; SceneScape [10] ensures geometric consistency through depth-aware generation; and Lumina-DiMOO [50] introduces discrete diffusion in image editing with impressive results. While these methods rely on specialized editing strategies, ViewMask-1-to-3 naturally achieves cross-view consistency through token-level modeling using discrete diffusion, offering a more unified and elegant solution to multi-view synthesis.
@@ 2.2 Discrete Diffusion Models
D3PM [2] established the foundational framework for discrete diffusion, extending traditional diffusion to discrete state spaces through structured denoising processes. DiffusionBERT [16] demonstrated that pre-trained BERT models could be adapted for text generation through discrete diffusion training. A significant breakthrough came with LLaDA [35], which showed that masked diffusion models can match autoregressive models like LLaMA3-8B in language generation while offering parallel decoding and bidirectional context modeling. LLaDA-V [54] extended this framework to multimodal understanding, demonstrating the potential of discrete diffusion in vision-language tasks. These advances have established discrete diffusion as a viable alternative to autoregressive approaches, for tasks requiring controllable and parallel generation.
@@ 2.3 Visual Tokenization
The effectiveness of discrete diffusion for visual tasks critically depends on visual tokenization quality. VQVAE [43] introduced discrete visual representations through vector quantization, laying the groundwork for discrete visual modeling. MaskGIT [6] applied masked token prediction to image generation using vector-quantized representations. MAGVIT-v2 [55] enabling high-quality discrete tokenization for both images and videos through lookup-free quantization and architectural improvements. The success of these tokenizers make discrete approaches increasingly competitive with continuous methods in visual generation tasks.
@@ 3.2 Model Design
We adopt MAGVIT-v2 [55] as our visual tokenizer due to its superior performance in discrete visual representation. For an input image $I_{i}$, the tokenizer produces a sequence of discrete visual tokens,
@@ 3.3 Multi-Stage Training Strategy
Stage1: Pretraining for Multimodal Alignment
@@
We construct a multimodal pretraining corpus of 1.2 million image–caption pairs using the images from ImageNet [9] images combined with captions provided by imagenet-1k-vl-enriched [24], to facilitate alignment between modalities.
@@
At this stage, the model is trained on a curated subset of Objaverse [8] and Habitat Synthetic Scene Dataset (HSSD) [21]. Following the data split provided by Objaverse++ [25], we removed samples labeled as Low Quality (No semantic meaning. Objects that annotators cannot identify or are corrupted), to ensure more stable training.
@@
We enhance Objaverse with Cap3D [33, 32], which offers detailed textual descriptions for 3D assets.
@@ 4 Experiments
Figure 4: Qualitative comparison on the 3D-Future dataset. Visual results on the 3D-Future dataset, where the bottom-right shows the reconstructed mesh, demonstrating that our ViewMask-1-to-3 produces more consistent and realistic novel views compared with recent methods.
@@
Our model was trained on a subset of Objaverse [8] and HSSD [21]. Following the data split provided by Objaverse++ [25], we removed samples labeled as Low Quality, to ensure more stable training. To enrich the textual supervision, we adopted the Cap3D [33, 32] annotations, which provide descriptive captions for 3D objects. In total, we used 180K 3D objects, each rendered as an 8-frame RGBA orbit sequence at a resolution of 256×256, with the elevation fixed at 30°. During training, one frame was randomly selected as the reference image for each object, resulting in a multi-view generation dataset.
@@
We evaluate the multi-view generation results from two perspectives. First, for the directly generated novel views, we adopt common image-level metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [44], and Learned Perceptual Image Patch Similarity (LPIPS) [56], to assess visual fidelity and perceptual quality. Then, we employ InstantMesh [51] to reconstruct 3D geometry from the generated views, and evaluate the resulting point clouds using Chamfer Distance (CD) and Intersection over Union (IoU) to quantify geometric consistency and accuracy.
@@ 4.1 Image to Multi-view Images
Method Architecture GSO 3D-FUTURE Rank
@@
Zero-1-to-3 [28] 2D Cont. Diff. 18.8219 0.8294 0.1659 17.0526 0.8163 0.1760 5.3
Zero-1-to-3 XL [28] 2D Cont. Diff. 19.6839 0.8381 0.1518 18.4702 0.8337 0.1485 3.0
ViVid-1-to-3 [22] 2D Cont. Diff. 19.7978 0.8566 0.1764 18.3241 0.8437 0.1682 3.5
Zero123++ [39] 2D Cont. Diff. 19.6373 0.8045 0.3550 23.5001 0.8527 0.1529 4.3
Epidiff [19] 2D Cont. Diff. 18.9917 0.8244 0.1667 17.4592 0.8108 0.1785 5.7
MV-Adapter [18] 2D Cont. Diff. 19.4673 0.7518 0.1503 19.3375 0.7217 0.4036 5.8
AR-1-to-3 [57] Video Cont. Diff. 13.5084 0.7376 0.3514 13.7724 0.7479 0.3034 8.5
TRELLIS [46] 3D Rectified Flow 14.4493 0.8058 0.3223 14.4604 0.7896 0.2990 7.2
@@
We compare our method against a comprehensive set of state-of-the-art diffusion-based novel view generation approaches, including Zero-1-to-3 [28], Zero-1-to-3 XL [28, 7], ViVid-1-to-3 [22], Zero123++ [39], MV-Adapter [18], EpiDiff [19], AR-1-to-3 [57], and TRELLIS [46].
@@
To quantitatively evaluate the zero-shot generalization capability of our model, we conduct experiments on two aspects: (1) image-level quality, measured on novel views generated from a single input image, and (2) 3D-level geometric accuracy, evaluated on 3D reconstructions. These experiments are evaluated on the Google Scanned Objects (GSO) dataset [50] and the 3D-FUTURE dataset [11].
The input views are selected with controlled offsets in azimuth and elevation relative to the canonical frontal view of each 3D object. For models that support arbitrary-angle rendering, such as Zero-1-to-3 and Zero-1-to-3 XL, the views are rendered at the same angles as those used by ViewMask-1-to-3 to ensure a fair comparison. We then evaluate PSNR, SSIM, and LPIPS following each model’s respective settings, and report the results in Tab. 1. Overall, our method, ViewMask-1-to-3, achieves the best combined performance across both datasets and attains the highest rank (1.7), demonstrating its superiority in multi-view generation.
To evaluate geometric consistency, we reconstruct 3D shapes from the generated multi-view images for each model. We then compute Chamfer Distance (CD) and IoU between the reconstructed shapes and the ground-truth meshes. The quantitative results are reported in Tab. 2, demonstrating that our method achieves lower CD and higher IoU, indicating more accurate 3D reconstruction and better cross-view consistency.